# FiDeLiS: Faithful Reasoning in Large Language Models for Knowledge Graph Question Answering

### v2

@@ Abstract
Large language models are often challenged by generating erroneous or ‘hallucinated’ responses, especially in complex reasoning tasks. To mitigate this, we propose a retrieval augmented reasoning method, FiDeLiS, which enhances knowledge graph question answering by anchoring responses to structured, verifiable reasoning paths. FiDeLiS uses a keyword-enhanced retrieval mechanism that fetches relevant entities and relations from a vector-based index of KGs to ensure high-recall retrieval. Once these entities and relations are retrieved, our method constructs candidate reasoning paths which are then refined using a stepwise beam search. This ensures that all the paths we create can be confidently linked back to KGs, ensuring they are accurate and reliable. A distinctive feature of our approach is its blend of natural language planning with beam search to optimize the selection of reasoning paths. Moreover, we redesign the way reasoning paths are scored by transforming this process into a deductive reasoning task, allowing the LLM to assess the validity of the paths through deductive reasoning rather than traditional logit-based scoring. This helps avoid misleading reasoning chains and reduces unnecessary computational demand. Extensive experiments demonstrate that our method, even as a training-free method which has lower computational costs and superior generality, outperforms established strong baselines across three datasets. The code of this paper will be released at https://anonymous.4open.science/r/FiDELIS-E7FC.
@@ 1 Introduction
The emergence and application of large language models (LLMs) (Brown et al., 2020; OpenAI, 2023; LLaMa-v3, 2024) have attracted widespread attention from researchers and the general public. They demonstrate remarkable reasoning capabilities, managing to solve complex reasoning problems through step-by-step thinking and planning (Wang et al., 2022b; Wei et al., 2022b). However, the reasoning of LLMs is not always reliable and may conflict with factual reality (Pan et al., 2024), limiting their application in areas requiring high reliability, such as healthcare (He et al., 2023) and science (Taylor et al., 2022).
Knowledge graphs (KGs) store high-quality knowledge in structured triplets, such as Wikidata (Vrandečić & Krötzsch, 2014), YAGO (Suchanek et al., 2007), and NELL (Carlson et al., 2010). KGs offer structured, explicit, and up-to-date factual knowledge, including domain-specific knowledge, providing a faithful knowledge source for reasoning. Moreover, each piece of information in KGs can be traced back to its source, providing context and provenance. This traceability not only aids in verifying the reliability of the information but also provides clear pathways of reasoning, making the interpretation process transparent. Due to its reliability and interpretability, it is considered as a promising method to improve the reliability of LLM reasoning. Therefore, several attempts have been conducted to integrate KGs with language models (Pan et al., 2024; Luo et al., 2024; Hu et al., 2023; Sun et al., 2023). Among them, knowledge graph question answering (KGQA) is a critical task to verify the effectiveness of incorporating the knowledge from KGs into reasoning models (He et al., 2021; Wang et al., 2023a; Yu et al., 2022). However, effectively integrating KGs into reasoning models for KGQA presents several challenges, including data sparsity, the complexity of query interpretation, and the need for advanced inference capabilities. More specifically, we focus on two main questions concerning the integration of KGs with reasoning models as follows:
(I) How to retrieve specific knowledge from KGs to allow more precise reasoning? Existing solutions include direct retrieval (Sun et al., 2019; Jiang et al., 2022) and semantic parsing (Sun et al., 2020; Gu & Su, 2022). Direct retrieval utilize the query to find relevant knowledge triplets within the KGs. However, these triplets sometimes lack comprehensive semantic information due to variations in KG schema, such as entities being represented as machine identifiers (MIDs), rather than descriptive labels. This lack of semantic richness can obscure the relevance of the retrieved information with the user query, and becomes especially tricky in multi-hop question answering, where seemingly unrelated intermediate triplets may be pivotal in deriving the correct answer. On the other hand, semantic parsing aim to transform user queries into executable structured queries (e.g., SPARQL), which are then run on KGs. However, these methods often grapple with issues related to the non-executability or incorrect execution of the generated queries (Yu et al., 2022), which undermines the relaibility of these methods.
(II) How to make the reasoning model understand and utilize the retrieved structured knowledge in KGs? Existing solutions include (1) finetuning LMs to generate relation paths grounded by KGs. These reasoning paths are then used for LLMs to conduct faithful reasoning and generate interpretable results (Luo et al., 2024; Yu et al., 2022); (2) leverage the reasoning model to iteratively retrieve and reason over subgraphs from the KG, deciding at each step to either answer the question, or to continue the searching step (Sun et al., 2023; Gu et al., 2023; Jiang et al., 2022). However, for the first approach, the reasoning steps generated by language models are not guaranteed to exist in the KG, especially when multiple consecutive steps are combined into a reasoning path. We conduct error analysis in Section 3 and show that only 67% of generated reasoning steps are valid, while the remaining 33% of reasoning steps either have a format error or do not exist in the KG. The latter solution faces the challenge of determining the optimal stopping point for the exploration process. For example, ToG (Sun et al., 2023) prompts LLMs to assess the adequacy of current reasoning paths for answer generation. However, assessing the adequacy of reasoning paths is itself a complex task that demands a deep understanding of the domain. This lack of clarity can easily result in premature stopping or excessive continuation, further complicating the decision-making process in LLMs.
To address these two challenges, we propose a retrieval augmented reasoning method, FiDeLiS. It is designed to enhance KGQA by anchoring responses to structured, verifiable reasoning paths. FiDeLiS is composed of two major components, Path-RAG, which retrieves chain of entities and relations from KGs in a effective manner (discussed in Section 2.1), and Deductive-verification Beam Search (DVBS), which conducts deductive-reasoning-based beam search to generate multiple reasoning paths leading to final answers (discussed in Section 2.2).
The Path-RAG module uses a keyword-enhanced retrieval mechanism that fetches relevant entities and relations from a vector-based index of KGs to ensure high-recall retrieval. Compared to a vanilla retriever which easily returns outputs that are not useful for finding the correct answer based on relevance to the query, we incorporate LLMs to generate an exhaustive list of keywords from the query to maximize coverage and ensure that no potential reasoning paths will be overlooked. Once these entities and relations are retrieved, the module constructs candidate reasoning steps which are then refined using a stepwise beam search discussed below.
Next, DVBS leverages the deductive reasoning capabilities of LLMs (Ling et al., 2024; Huang & Chang, 2023) as a clear-defined criterion for automatically directing the beam search process step by step to create the complete reasoning paths. A key characteristic of DVBS is its integration of natural language planning (Song et al., 2023; Zhou et al., 2022) and beam search. This combination enhances the process of choosing the best reasoning paths by providing additional hints for decision making in LLMs. In addition, we redesign how reasoning paths are evaluated by converting the scoring process into a deductive reasoning task. This allows the LLM to validate the reasoning paths through deductive verification rather than traditional logit-based scoring. This proposed deductive verification serves as presie indicators for when to cease further reasoning, thus avoiding misleading reasoning chains and reducing the computational resources required. Overall, our main contributions include:
  • We propose a retrieval augmented reasoning method, which enhances knowledge graph question answering by anchoring responses to structured, verifiable reasoning paths grounded by KGs.
  • We propose a step-wise keyword retrieval method that enhances the recall of relevant intermediate knowledge from KGs. This ensures that all the paths we create can be confidently linked back to KGs, ensuring that they are faithful and reliable.
  • We propose deductive verification as precise indicators for when to cease further reasoning, thus avoiding misleading the chains of reasoning and reducing unnecessary computation required.
  • Extensive experiments show that our method, as a training-free method with lower computational cost and better generality, outperforms existing strong baselines in three datasets.
@@ 2 Method
[…]
Overview. The framework of FiDeLiS is as shown in Figure 1. Given a query, it first retrieves the reasoning step candidates from KGs (as discussed in Section 2.1). To ensure high recall, our method first generates a list of keywords from the query. And then based on the generated keywords, it retrieves entities and relations from an index (where the entities and relations are also embedded by vector representation obtained from an LM). Next, following the retrieved entities, n-hop reasoning step candidates are created and scored based on semantics similarity and neighborhoods aggregation. Then, a deductive verification-based beam search (discussed in Section 2.2) is employed to generate the top-k reasoning paths leading to the answer. First, a natural language plan is generated from the query to provide more hints for LLMs decision making and then a beam search procedure is conducted to score each reasoning step. We redesign the scoring function by converting the process into a deductive reasoning task. This allows the LLM to validate connections between paths through simple verification rather than traditional logit-based scoring. This change helps prevent easily misleading reasoning chains and reduces the computational resources required.
Figure 1: Overview of FiDeLiS. The goal of FiDeLiS is to extend the multiple-step reasoning paths $T$ grounded by KG. The figure shows given a query, Path-RAG first retrieve entities and relations that are relevant to the query, and construct the reasoning path candidates. Then DVBS constructs reasoning paths by iterative extending these path candidates using beam search. Specifically, at timestep $t$, DVBS leverage LLMs to choose the top-$k$ (here we set $k=2$ as an example) reasoning steps and decide whether to continue the next search step or cease the reasoning path extension based on deductive verification. At last, FiDeLiS return the top-$k$ reasoning paths as the final output leading to the answers of the query.
@@ 2.1 Path-RAG: Reasoning Path Retrieval-Augmented Generation
Path-RAG aims to retrieve reasoning path candidates from KGs and comprises three main steps: initialization, retrieval, reasoning step candidates construction, as depicted in Figure 1. The implementation details of each step are elaborated in the following paragraphs.
Initialization. We initiate the Path-RAG by generating embeddings for entities (nodes) and relations (edges) using a pre-trained language model (LM). We begin by extracting entities ($e_{i}$) and relations ($r_{i}$) from the KG, denoted as $\mathcal{E}$ and $\mathcal{R}$, respectively. Each entity and relation is then encoded using a pre-trained language model such as SentenceBert (Reimers & Gurevych, 2019) or E5 (Wang et al., 2022a), yielding dense vector representations:
$z(e^{i})=\operatorname{LM}(e_{i})\in\mathbb{R}^{d},\ z(r^{i})=\operatorname{LM}(r_{i})\in\mathbb{R}^{d}$ (1)
where $d$ denotes the dimension of the output vector. These embeddings are stored in a nearest neighbor data structure, which facilitates quick retrieval of similar entities or relations based on their vector representations.
Retrieval. To retrieve the entities and relations that are relevant to the user query, the embeddings of them are utilized to populate a nearest neighbor index. Specifically, for a given query $q$, we use LLM to generate an exhaustive list of keywords or relation names referred to as $\mathcal{K}$ (the prompt is given in Section F). These keywords are generated to maximize the coverage of potential reasoning steps essential for answering the query. The list is derived as $\mathcal{K}=\text{LM}(\text{prompt}_{p},q)$. Then each keyword $\in\mathcal{K}$ is concatenated into a string denoted as $K$ and then encoded into the same latent space by the same LM used in initialization step to obtain $z(K)=\text{LM}(K)\in\mathbb{R}^{d}$.
To find the entities and relations that best match the keyword embeddings, we compute the cosine similarity $\cos(\cdot,\cdot)$ between $z(K)$ and the embeddings of entities ($z(e)$) and relations ($z(r)$) in the KG. We then select the top-$m$ elements that exhibit the highest similarity:
$\mathcal{E}_{m}=\operatorname{argtopm}_{i\in\mathcal{E}}\cos\left(z(K),z(e)\right),\mathcal{R}_{m}=\operatorname{argtopm}_{i\in\mathcal{R}}\cos\left(z(K),z(r)\right)$ (2)
Each entity and relation in the retrieved set is then assigned a score, reflective of its similarity to the keyword embedding as $S_{\text{ent}}(e)=\cos\left(z(K),z(e)\right)$ and $S_{\text{rel}}(r)=\cos\left(z(k),z(r)\right)$.
Reasoning Step Candidates Construction. After the initial retrieval process, which runs once at the start of the algorithm, we iteratively construct reasoning step candidates to extend the reasoning paths based based on entities $\mathcal{E}_{m}$ and relations $\mathcal{R}_{m}$ that are relevant to the given query. To guide the selection of these candidates, we propose another scoring function based on the derived relevance score $S_{\text{ent}}(e)$ and $S_{\text{rel}}(r)$. Our scoring function leverages triplets to represent the connectivity of KGs and assess whether extending a path in certain directions might lead to negative outcomes over multiple steps by incorporating next-hop neighbor’s information.
$S((r,e))=S_{\text{rel}}(r)+S_{\text{ent}}(e)+\alpha\max_{\forall(r_{j},e_{j})\in N(e)}\left(S_{\text{rel}}(r_{j})+S_{\text{ent}}(e_{j})\right)$ (3)
Here, $(r,e)$ refers to the reasoning step defined as a relation-entity pair. $S_{\text{rel}}(r_{j})$ and $S_{\text{ent}}(e_{j})$ are the similarity scores for relations and entities at the next step. The component $N(e)$ corresponds to entities reachable from $e$ within one hop. The factor $\alpha$ is used to balance short-term outcomes and long-term potential in reasoning paths: a higher $\alpha$ prioritizes paths with long-term benefits, even if they seem sub-optimal initially, while a lower $\alpha$ emphasizes immediate steps, potentially overlooking future impacts. Details of Path-RAG are provided in Algorithm 1.
2.2 DVBS: Deductive-Verification Guided Beam Search
DVBS is designed to prompt LLMs to iteratively execute beam search on the reasoning step candidates (provided by Path-RAG) to find the top-$k$ most promising reasoning paths that lead to the answer to the user query. At each timestep, we leverage the LLMs to choose the top-$k$ reasoning steps and decide whether to continue extending the reasoning paths or cease the extension process. It comprises three main steps: planning, beam-search, deductive-verification, as depicted in Figure 1. The implementation details of each step are elaborated in the following paragraphs.
Plan-and-Solve. Inspired by the recent works regarding planning capabilities of LLMs (Song et al., 2023; Zhou et al., 2022), we prompt LLM to generate the planning steps for answering the user query, denoted as $w$. This step aims to provide more hints for subsequent LLM decision making process. The detailed prompt can be found in Section F.
Beam Search. To enable multi-step reasoning, a reasoning path of $T$ steps is sequentially generated through several time steps as $\left[s^{1},s^{2},\ldots,s^{t},s^{T}\right]=s^{1:T}$, where $s^{t}$ represents the reasoning step at timestamp $t$. We constrain each reasoning step should within a set of candidates $\mathcal{S}^{t}$ to control the computation efficiency. We construct the reasoning step candidates $\mathcal{S}^{t}$ based on the scoring function defined in Eq. 3 by selecting the top-$m$ reasoning steps with the highest scores. At each timestamp $t$, we leverage the LLMs to select the top-$k$ reasoning steps from the reasoning step candidates $\mathcal{S}^{t}$. The beam search process at timestamp $t$ is modeled as follows:
$\mathcal{H}_{t}=\text{Top}_{k}\left\{h\oplus\text{LM}(s^{t}|q,s^{1:t-1},w):h\in\mathcal{H}_{t-1},s^{t}\in\mathcal{S}^{t}\right\}$ (4)
where $\mathcal{H}_{t-1}$ denote the reasoning path up to the previous timestamp $t-1$ and $s^{t}$ refers to the current reasoning step. $\mathcal{S}^{t}$ refers to the set of possible next-step candidates retrieved by the Path-RAG at timestamp $t$. $\text{LM}(s^{t}|q,s^{1:t-1},w)$ refers to the language models prediction for the next step $s^{t}$ given the previous sequence $s^{1:t-1}$, the query $q$, and the planning context $w$. The operator $\oplus$ appends $s^{t}$ to the current path $h$, and $\text{Top}_{k}$ selects the top-$k$ reasoning paths (with $k$ controlling the beam search width, where a larger $k$ typically yield better performance (see the analysis in Figure 2)).
Deductive Verification. To terminate the reasoning path extension process at the right point, we propose to leverage the deductive reasoning capabilities of LLMs (Ling et al., 2024; Huang & Chang, 2023) as a criterion $C(x,s^{t},s^{1:t-1})\in\left\{0,1\right\}$ to automatically guide the decision making of LLMs in Equation 4. Specially, we first leverage LLMs to convert the user query to a declarative statement $q^{\prime}$ (as shown in Appendix F.6) and then use LLMs to judge whether it can be deduced from current reasoning step $s^{t}$ and the previous reasoning steps $s^{1:t-1}$. We utilize the same backend LLM with prompt referred in Section F and define the deductive-verification criteria as follows:
$C(q^{\prime},s^{t},s^{1:t-1})=\begin{cases}1,&\text{if }q^{\prime}\text{ can be deduced from }s^{t}\text{ and }s^{1:t-1},\\ 0,&\text{otherwise.}\end{cases}$ (5)
The overall goal of DVBS model can be represented as:
$\mathcal{H}_{t}=\text{Top}_{B}\left\{h\oplus\text{LM}(s^{t}|q,s^{1:t-1},w):h\in\mathcal{H}_{t-1},s^{t}\in\mathcal{S}\text{ and }C(q^{\prime},s^{t},s^{1:t-1})=1\right\}$ (6)
This criterion is an essential component of our method for leveraging LLMs to iteratively validate each step of reasoning, ensuring that each step logically follows from the preceding steps and aligns with the original query. The function effectively signals when to terminate further reasoning, enhancing accuracy and minimizing unnecessary computational efforts.
3 Experiments
During the experiment section, we aim to answer the following research questions: RQ1: How does FiDeLiS perform compare with existing baselines on KGQA tasks? RQ2: How does the design of each individual component contribute to the overall performance of FiDeLiS? RQ3: How robustness of FiDeLiS perform under varying conditions? RQ4: How about the efficiency of FiDeLiS? All experiment settings are detailed in Appendix D due to page constraints.

### v3

@@ Abstract
Large language models (LLMs) are often challenged by generating erroneous or hallucinated responses, especially in complex reasoning tasks. Leveraging knowledge graphs (KGs) as external knowledge sources has emerged as a viable solution. However, existing KG-enhanced methods, either retrieval-based or agent-based, encounter difficulties in accurately retrieving knowledge and efficiently traversing KGs at scale. In this paper, we propose a unified framework, FiDeLiS, designed to improve the factuality of LLM responses by anchoring answers to verifiable reasoning steps retrieved from a KG. To achieve this, we leverage step-wise beam search with a deductive scoring function, allowing the LLM to validate each reasoning step and halt the search once the question is deducible. In addition, our Path-rag module pre-selects a smaller candidate set for each beam search step, reducing computational costs by narrowing the search space. Extensive experiments show that our training-free and efficient approach outperforms strong baselines, enhancing both factuality and interpretability. Code is released at https://anonymous.4open.science/r/FiDELIS-E7FC.
FiDeLiS: Faithful Reasoning in Large Language Models for Knowledge Graph Question Answering
Yuan Sui1, Yufei He1, Nian Liu1, Xiaoxin He1, Kun Wang1,2, Bryan Hooi1 1National University of Singapore 2University of Science and Technology of China {yuansui, yufei.he, nianliu, xiaoxin, bhooi}@comp.nus.edu.sg, wk520529@mail.ustc.edu.cn
@@ 1 Introduction
Large language models (LLMs) have shown impressive reasoning capabilities in tackling complex tasks Yu et al. (2024). However, the reasoning of LLMs is not always reliable and can be prone to generating outputs that are either inconsistent with real-world facts Xu et al. (2024); Huang et al. (2025) or show flawed reasoning process Li et al. (2024); Sui et al. (2024), which greatly undermine the reliability of LLMs in real-world applications Kung et al. (2023); Zhang et al. (2024).
Figure 1: Challenges for existing KG-enhanced methods: How to balance faithfulness and efficiency?
To address this issue, leveraging knowledge graphs (KGs) as external knowledge sources has emerged as a viable solution Sun et al. (2023); Ma et al. (2024); Luo et al. (2024a). Unlike traditional retrieval-augmented generation (RAG) that relies on web pages or documents Liu et al. (2024); Qian et al. (2024); Bayarri-Planas et al. (2024), KGs represent information in a structured and interconnected format, where each fact is stored as entities and relations. This format supports explicit, traceable reasoning processes Pan et al. (2023) and facilitates multi-hop reasoning through graph traversal. Moreover, each fact in a KG can be traced back to its source Sui et al. (2024); Agrawal et al. (2024), providing both context and original details, which further enhances the information authenticity and reliability of the reasoning processes.
Existing KG-enhanced LLM reasoning methods face notable challenges and can be roughly categorized into two primary approaches: retrieval-based and agent-based paradigms Luo et al. (2024b). Retrieval-based methods Wang et al. (2023); Luo et al. (2024a); Baek et al. (2023) retrieve relevant KG facts to support LLM reasoning by either prompting Baek et al. (2023) or fine-tuning LLMs to learn the underlying structure of KG Luo et al. (2024a, b). These methods often suffer from incomplete or imprecise information extraction due to a lack of contextual understanding or an inability to fully capture the graph structure Luo et al. (2024b).Our error analysis of a strong retrieval-based method (i.e., Luo et al. (2024a)) in §4.3 reveals that only 67% of the generated reasoning steps are valid, with 33% containing format errors or referencing non-existent KG facts. In contrast, agent-based methods Sun et al. (2023); Ma et al. (2024) treat LLMs as interactive agents that explore KGs iteratively to construct reasoning paths and generate answers. While this approach by nature can enhance reasoning accuracy, it is computationally expensive, resulting in high latency and scalability limitations. As illustrated in Figure 1, balancing faithfulness and efficiency remains a critical challenge for existing KG-enhanced reasoning methods.
To this end, we propose FiDeLiS, a unified framework designed to improve the factual accuracy and reasoning efficiency of LLMs on KGQA task. FiDeLiS anchors LLM responses to verifiable reasoning steps derived from a KG by employing two core components: (1) Deductive-Verification Beam Search (DVBS) which systematically constructs and validates reasoning paths step-by-step, ensuring logical consistency and factual correctness (discussed in §3.2). This module also prevents premature reasoning termination and incorrect path extension to ensure the validity of the generated reasoning paths. (2) Path-RAG, a retrieval-augmented mechanism that pre-selects a constrained set of candidate entities and relations for each step to mitigate computational inefficiencies. It combines semantic similarity measures with graph-based connectivity analysis to optimize the search space, significantly reducing latency without sacrificing recall or accuracy (discussed in §3.1). Extensive experiments show that our method outperform strong baselines in both accuracy and efficiency, offering a scalable, training-free solution for KG-enhanced LLM reasoning. Overall, our main contributions include:
  • We propose FiDeLiS, a unified framework designed to improve the factual accuracy of LLMs by grounding reasoning paths in structured KG efficiently.
  • We enable efficient, verifiable reasoning by deductively validating reasoning steps and narrowing the search space with high-quality retrieval mechanism.
  • FiDeLiS performs robustly across different experiments without the need for model fine-tuning, demonstrating adaptability and scalability with improved performance on multiple benchmarks.
  • By anchoring responses in verifiable reasoning paths, FiDeLiS enhances interpretability, enabling users to verify and understand each reasoning step.
2 Preliminary
Notation. To facilitate the demonstration of our method, we define the necessary notation below:
  • Definition 1. A reasoning step is a pair $(r,e)$, where $r$ is the relation and $e$ is the corresponding entity.
  • Definition 2. A reasoning path $\mathcal{P}$ is a pair $(s,\mathcal{T})$, where $s$ is the starting entity for the reasoning path, and $\mathcal{T}$ is a sequence of reasoning steps $\mathcal{T}=\left\{t_{1},\ldots,t_{n}\right\}$ and $t_{k}=(r_{k},e_{k})$ denotes the $k$-th reasoning step in the path and $n$ denotes the length of the path.
  • Definition 3. The next-hop candidates given path $\mathcal{P}$, denoted $\mathcal{N}_{1}(e_{n})$, is defined as the 1-hop neighborhood of $e_{n}$, the last node in the reasoning path $\mathcal{P}$.
[…]
Task definition. In this work, we focus on the task of knowledge graph-based question answering (KGQA), a common reasoning task involving KGs. It is defined as: given a user query $q$ and a KG $\mathcal{G}=\{(e,r,e^{\prime})\mid e,e^{\prime}\in\mathcal{E},r\in\mathcal{R}\}$, where $\mathcal{E}$ and $\mathcal{R}$ denote the set of entities and relations in KG, the task aims to design a function $f$ to predict answers $a\in\mathcal{A}_{q}$ conditioned on $q$ and $\mathcal{G}$. Following existing KG-enhanced LLMs methods Sun et al. (2023); Ma et al. (2024), the function $f$ can be generally expressed as finding valid reasoning path(s) $\mathcal{P}$ on KGs that connects the entities mentioned in the query and the answer as: $P(a|q,\mathcal{G})=\sum_{\mathcal{P}}P_{\theta}(a|q,\mathcal{P})P_{\phi}(\mathcal{P}|q,\mathcal{G})$, where $P_{\theta}(a|q,\mathcal{P})$ denotes the probability of generating answer $a$ conditioned on $q$ and reasoning path(s) $\mathcal{P}$ by a function parameterized by $\theta$, and $P_{\phi}(\mathcal{P}|q,\mathcal{G})$ denotes the probability of discovering reasoning path(s) $\mathcal{P}$ by a function parameterized by $\phi$. As reasoning path $\mathcal{P}$ is defined as a sequence of reasoning steps, we factorize the reasoning path probability using the chain rule as Eq 1:
$P(a|q,\mathcal{G})=\sum_{\mathcal{P}}P_{\theta}(a|q,\mathcal{P})\prod_{k=1}^{n}P_{\phi}(t_{k}|q,t_{<k},\mathcal{G})$ (1)
To acquire valid reasoning paths, most prior studies follow the retrieval-based Li et al. (2023); Luo et al. (2024a) or agent-based Sun et al. (2023) paradigm. As indicated in Luo et al. (2024b), retrieval-based methods rely on precise additional retrievers, while agent-based methods are computationally intensive and lead to high latency. To address these issues, we propose our method, FiDeLiS, to enable both efficient and faithful reasoning over KGs.
Figure 2: An illustration of FiDeLiS. Top: The workflow of Path-RAG. An LLM first extracts key terms and generates dense embeddings that feed into the Path-RAG module. Path-RAG rapidly retrieves relevant entities and relations from a pre-embedded KG and constructs candidate reasoning steps by combining semantic similarity with graph connectivity. Bottom: The workflow of DVBS. Then, the DVBS module uses an LLM-generated planning outline to guide a beam search that builds reasoning paths step-by-step over candidates constructed by Path-RAG, with deductive verification ensuring each step logically follows the previous steps and support the user question.
@@ 3 Method
Motivated by the insight that integrating KGs with LLMs can mitigate hallucinations and enable verifiable reasoning, we propose FiDeLiS to improve the factuality of LLM responses by anchoring answers to verifiable reasoning steps retrieved from a KG. The overall framework of FiDeLiS is illustrated in Figure 2, which consists of two main components: (1) Reasoning Path Retrieval-Augmented Generation (Path-RAG, Algorithm 1) and (1) Deductive-verification Beam Search (DVBS, Algorithm 2).
Given a complex question $q$, we first use an LLM to extract key terms from $q$ and generate dense embeddings that capture the question’s core concepts. These embeddings are input into Path-RAG module, which rapidly retrieves relevant entities and relations from a pre-embedded KG to select a smaller candidate sets for further beam search step, addressing the latency and computational burden of traditional agent-based methods. Path-RAG then constructs candidate reasoning steps by combining immediate semantic similarity with the structural connectivity of the graph, overcoming the dependence on highly precise retrievers in standard retrieval-based approaches. Next, the DVBS module employs an LLM-generated planning outline to guide a beam search that builds reasoning paths step-by-step. At each step, deductive verification checks that the accumulated reasoning steps logically supports the user question, ensuring the final reasoning path is both verifiable and accurate.
@@ 3.1 Path-RAG: Reasoning Path Retrieval-Augmented Generation
Previous agent-based methods Ma et al. (2024); Sun et al. (2023) treat LLMs as agents that iteratively interact with KGs to find reasoning paths and answer, which necessitate multiple rounds of interaction between agents and KGs and lead to high computational costs and latency. We instead propose a module, Path-RAG which iteratively pre-select a smaller candidate set to reduces the search space for exploring the potential reasoning paths from KGs. It consists of three steps and we detail the workflow as follows:
Initialization.
We initiate the Path-RAG by encoding each entity $e_{i}\in\mathcal{E}$, and relation $r_{i}\in\mathcal{R}$ in the KG using a pre-trained language model (LM), which produces dense vectors $z(e^{i})=\mathrm{LM}(e_{i})\in\mathbb{R}^{d}$ and $z(r^{i})=\mathrm{LM}(r_{i})\in\mathbb{R}^{d}$, where $d$ denotes the embedding dimension. These embeddings are stored in a nearest neighbor structure to facilitate rapid similarity search.
Keyword-Driven Retrieval.
We then populate a nearest neighbor index to retrieve relevant entities and relations for the user query. We first use an LLM to analysis the user query and generate exhaustive keywords/relation names that could be useful for finding the reasoning path to answer the query (See the prompt in §C.1). This step is designed to maximize coverage of potential reasoning steps, ensuring that no potential reasoning paths are overlooked during the retrieval process. The extracted keywords are then encoded using the same LM in initialization, yielding $z(K)=\mathrm{LM}(K)\in\mathbb{R}^{d}$. We subsequently compute the cosine similarity between $z(K)$ and the pre-stored embeddings, retrieving the top-$m$ entities and relations: $\mathcal{E}_{m}=\operatorname{argtopm}_{i\in|\mathcal{E}|}\cos\left(z(K),z(e_{i})\right)$ and $\mathcal{R}_{m}=\operatorname{argtopm}_{i\in|\mathcal{R}|}\cos\left(z(K),z(r_{i})\right)$.
Reasoning Step Candidates Construction.
Next, we construct candidate reasoning steps defined in §2 using the retrieved candidate entities $\mathcal{E}_{m}$ and relations $\mathcal{R}_{m}$. To guide the selection of potential candidate, we propose a scoring function that combines semantic similarity with the KG’s structural connectivity. First we define the base score function $S_{0}$ that captures only the semantic alignment of the candidate with the query as: $S_{0}((r,e))=S_{\mathrm{rel}}(r)+S_{\mathrm{ent}}(e)$, where $S_{\mathrm{ent}}(e)$ and $S_{\mathrm{rel}}(r)$ represents the cosine similarity between the entity/relation to the query respectfully. To account for the KG’s structural connectivity, i.e., the potential for a candidate to lead to fruitful next steps, we incorporate information from the next-hop candidates and define the overall scoring function as Eq 2:
$S((r,e))=S_{0}((r,e))+\alpha\max_{\forall(r_{j},e_{j})\in N(e)}S_{0}((r_{j},e_{j}))$ (2)
Where $N(e)$ denotes the set of candidate relation-entity pairs reachable from entity $e$ within one hop in the KG. $\alpha$ is a hyper-parameter that balances the immediate semantic relevance (captured by $S_{0}((r,e))$ with the candidate’s potential for future connectivity (captured by the maximum next-hop score). A higher $\alpha$ favors candidates with long-term benefits, even if they seem sub-optimal initially, while a lower $\alpha$ emphasizes immediate rewards, potentially overlooking future impacts. We verify the effectiveness of this new scoring function in Table 3 and append the tuning results of hyper-parameter $\alpha$ in Figure 5.
3.2 Deductive-Verification Beam Search
The objective of DVBS is two-fold: (1) to provide a step-wise beam search for exploring verifiable reasoning paths from KG based on candidates constructed by Path-RAG (§3.1), and (2) to verify each reasoning step based on deductive reasoning Ling et al. (2023) to ensure each step logically follows the previous steps and supports the user query. Compared with existing methods like Sun et al. (2023) and Ma et al. (2024), while we both consider treat LLMs as agents that iteratively interact with KGs to find reasoning paths and answers, our method leverage deductive reasoning to ensure each reasoning steps are logically connected and only halt the search if the question can be deduced based on the reasoning paths. Based on our robustness analysis in §4.3, DVBS demonstrate higher ratio of valid reasoning paths and can prevent issues of either premature stopping Huang et al. (2017) or excessive continuation of reasoning path extension. The DVBS consists of three steps and we detail the workflow as follows:
Plan Generation.
Inspired by the recent works regarding planning capabilities of LLMs Zhang et al. (2023); Kagaya et al. (2024), we prompt an LLM to generate the planning steps for answering the user query, denoted as $w$. This step is designed to provide more hints for subsequent LLM decision making process. Even this step is more like an engineering trick, we find that it may unlock some of the capabilities of LLM to do “higher-order” thinking. By including more hints in the prompt, the LLM tends to make more accurate and deterministic decisions during beam search, thus improving the quality of the traversed reasoning paths. (See Table 2 for ablation analysis).